### v1

@@ 1.1 Predicting and observing stellar occultations from space
In June 2020, the first-ever predicted stellar occultation observation using a space telescope was obtained with the Characterising ExOPlanets Satellite (CHEOPS) [36]. This observation demonstrated the feasibility of predicting and observing stellar occultations caused by distant solar system minor bodies using space telescopes: these facilities are unaffected by weather conditions and can extend the range of observations beyond the Earth’s surface. The use of space telescopes also facilitates the detection of material surrounding small bodies, as it enables photometric accuracy comparable to that achieved by ground-based telescopes with much larger apertures. We also have access to stellar occultations of fainter stars than from the ground, multiplying the number of possible events.
@@
JWST is orbiting close to the second Sun-Earth Lagrange point (L2) [37, 38]. L2 is an unstable equilibrium point, meaning constant manoeuvres must be made to keep the telescope in orbit around it [39]: such manoeuvres are usually made 2–3 times a month. Consequently, the predicted orbital path of JWST needs to be frequently updated (J. Giorgini, personal communication, June 22, 2022). For all the above reasons and due to the complex motion of the space telescope around L2, predicting stellar occultations by distant minor bodies visible from JWST is a very challenging task. Extended Data Figures shows the difference between two consecutive releases of the JWST ephemeris, obtained on September 27, 2022. In just under six months from the initial epoch, the difference between the two releases reached approximately 10,000 km, which is larger than the Earth’s radius. In the absence of formal positional uncertainty, the plot illustrates the trustworthiness of a stellar occultation prediction using JWST far from the initial epoch.
@@
To predict stellar occultations that JWST could potentially observe, we used the Stellar Occultation Reduction and Analysis (SORA) software package [40] combined with the Gaia DR3 catalogue [41, 42] and the NIMA ephemeris [43] for the solar system bodies. The strategy we developed to run the predictions is as follows. We selected the 40 largest TNOs/Centaurs (36 TNOs and 4 Centaurs), then ran predictions for these 40 targets every week for Gaia DR3 stars up to a G-magnitude of 20. This frequency of re-running the predictions was necessary due to the station-keeping manoeuvres applied to JWST, which changed the position (ephemeris) of the space telescope. Each week, we reviewed the updated predictions to identify favourable events, refined these predictions with new astrometric data obtained from the ground, and reran the predictions for those objects. Finally, when we were confident in the prediction, we activated the target of opportunity (ToO) to observe the occultation from JWST. Since this was a non-disruptive ToO, it had to be activated at least 14 days in advance of the event.
@@
Using the described strategy, we identified the occultation by the ringed Centaur Chariklo on 18 October 2022 (UT), with a low relative velocity (2.5 km/s) and a star bright enough to be observed by JWST (K = 15.7 mag) as the most promising event. We first identified this occultation on 19 August 2022 (blue line in Extended Data Figures). Then we reran the prediction every week to account for changes in JWST’s orbit caused by station-keeping manoeuvres. Our most recent prediction was made on the day of the occultation, just a few hours before the event (green line in Extended Data Figures). In just about two months, the path of the star behind Chariklo, as seen by JWST, moved 110 km, enough to miss the occultation by the main body. Furthermore, the instant of occultation also shifted by 1.5 minutes.
@@ 1.2 Observations
We used the James Webb Space Telescope (JWST) to observe a stellar occultation by Chariklo’s rings, marking the first time such an event was both predicted and successfully observed by JWST. It also represents the first stellar occultation by a small Solar System body ever recorded from beyond Earth’s orbit, with the only comparable case being a previous Quaoar occultation observed by the Earth-orbiting CHEOPS satellite. This observation was both scientifically valuable and technically demanding. It was conducted under JWST’s Solar System Guaranteed Time Observations (GTO) program #1271, ToO TNOs: Unveiling the Kuiper belt by stellar occultations (PI: Pablo Santos-Sanz) [44].
@@ 1.3 Photometry: occultation light curves
Fluxes from Chariklo and the target star were also measured separately in both filters, using images acquired before the occultation when the two sources were spatially resolved (Extended Data Figures, middle Panel). These measurements were used to calibrate the relative contributions of Chariklo and the star to the combined flux, a crucial step in deriving the ring parameters discussed in Section 1.4. The fractional flux contribution of Chariklo (relative to the combined Chariklo + star flux) was 0.2780 for F150W2 and 0.2662 for F322W2.
@@ 1.4 Analysis of the rings
In the obtained light curves, the main, densest, and most opaque ring of Chariklo is clearly visible. This ring is called C1R, following the original notation [7]. The C1R occultation light curve displays obvious diffraction spike features immediately before and after the ingress and egress, respectively. These spikes provide strong evidence that the ring has sharp edges and is tightly confined, similar to the rings of Uranus [23, 24]. However, the dynamical mechanism responsible for this confinement is still unknown and is currently under discussion [16, 45, 46, 33].
The other ring (referred to as C2R) is surprisingly not clearly detected in the F150W2 or F322W2 light curves, as shown in Extended Data Figures. Since a ring partially occults the starlight, we fit an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction, finite bandwidth, exposure time, and the Gaia DR3 stellar diameter. We use the procedures within the SORA package [40]. From this analysis, we obtained the time when the star disappeared and reappeared behind Chariklo’s rings ( and ) and the observed apparent ring opacity (), which measures the fractional drop of stellar flux.
Knowing the reconstructed JWST position at the time of the occultation, we were able to propagate the instants of disappearance and reappearance of the star into the tangent plane and determine two-dimensional coordinates (, ) of the rings’ positions relative to Chariklo’s expected centre. This projection was executed using the SORA package, and its result is shown in Figure 1. Assuming Chariklo’s parameters (body and rings) as published by [15], we fitted the centre position of Chariklo as viewed by JWST at the mid-point of the occultation (18 October 2022 at 13:37:42.18 UT), which resulted in
That meant an appulse with the closest approach of 132.2 km, in contrast to Chariklo’s volume-equivalent radius of 124.8 km [15].
Initially, the parameters of both rings (C1R and C2R) were obtained in the sky plane simultaneously, just before and after the closest approach. We used the formalism presented by [14] to project these parameters to the ring plane so we could obtain the relevant parameters, such as the ring radial width (), the normal opacity (), and the normal optical depth ().
@@
alternatively, if we assume a poly layer ring, is defined as [47]
where is the ring opening angle, and (resp. ) corresponds to a pole-on (resp. edge-on) viewing. On the other hand, considering that the Airy scale is larger than the width of the ring as seen in the sky plane [48], a factor is needed to compute the fraction of lost light; in that case
@@
The integrals of and over the radial width () of the ring define the equivalent width () and the equivalent depth () of the profile. These values ( and ) are proportional to the amount of material present in the profiles for each case [14]. The ring parameters are obtained in Table 1. We highlight that the detection of C2R is 1 significant in the F150W2 data and cannot be detected in the F322W2 data.
@@ 1.5 Sampling a ring with azimuthally variable structure
In the case of a ring with an azimuthally variable structure, we may observe different normalized opacities () or equivalent with () values at the 1 and 2 contacts during an occultation event, as well as variations between multiple occultations. If the ring opacity is assumed to be wavelength-independent, the observed decrease/increase in opacity and/or in during the 2022 JWST events—relative to previous occultations—could be explained by the possibility that both ingress and egress sampled low/high-density segments of the ring. To estimate the likelihood of such an occurrence, we performed a simple Monte Carlo simulation. We modelled an azimuthally variable ring, with the material distribution given by or , where is the azimuthal angle and is a small integer. Random chords drawn through this ring yield a distribution of ‘observed’ or values.
In the real occultation measurements of the C2R ring, we obtain in seven of the occultations (see Extended Data Tables), with significantly lower values observed only during the JWST event. For simplicity, we assume that the minimum of along the ring is zero, although a constant positive offset may exist. Therefore, the probability of randomly sampling a pair of such low values in just one of eight events, as calculated from the simulation, represents an upper limit, with a probability of obtained in our simulation. In a similar way, we obtain a probability of p 0.004, assuming that we sample the ring close to the density maxima in the JWST measurements, and below these values in the pre-JWST cases (Extended Data Tables). While azimuthal fragmentation is plausible—especially given the existence of fragmented rings in the trans-Neptunian region, e.g. in the case of Quaoar [36, 22]—these very low probabilities suggests that the observed JWST opacity drop (C2R) or increase (C1R) is unlikely to be solely due to this effect. The results have been found to be independent of the parameter for 3. Here, we assumed that the azimuthal structures of the two rings are independent.
@@ 1.6 Spreading timescale of Chariklo’s C2R Ring
We compute the characteristic spreading timescale of Chariklo’s second ring (C2R) using the standard expression for a narrow, self-gravitating ring in resonance [49]:
@@
We adopted ellipsoidal semi-axes from Morgado et al. (2021)[15]: km, km and km and a mean bulk density of .
@@
To account for the uncertainty in radial width, we uniformly sampled over the KDE-inferred interval [0.40, 3.25] km. A Monte Carlo simulation with realisations yielded a median spreading timescale of s (0.36 yr), with a 1 confidence interval between s (0.23 yr) and s (0.77 yr).
@@ 1.7 Radiative Transfer Modeling
After the striking marginal (at F150W2) and non-detection (at F322W2) of the C2R in the JWST filters, wavelength-dependent observables induced by micron- and submicron-sized grains became a real concern in explaining the measured discrepancies against previous occultations. Here, we perform a thorough analysis of the compatibility of dusty ring models for both C1R and C2R. Our main goals are to investigate which material in our models can replicate these observations and what we can say about the possible presence or absence of water ice in the rings. To explore whether a potential wavelength-dependent effect is responsible for the observed changes in the measured opacities, we conducted a radiative transfer analysis, assuming that the ring did not change during the observation period (2013-2022) and that all observational chords sampled the same amount of material, after correcting for the opening angle change due to Chariklo’s orbital motion around the Sun. We used a slab-like, two-component model, where we mixed materials to match their common optical properties with the measured occultation normal opacities () or, equivalently, their normal optical depths (). In our model, the optical depth is obtained as the sum of the optical depths of two materials (two different grain sizes and/or compositions):
@@
where is the slab’s length (the ring’s thickness). A single-component model is obtained if is chosen. To calculate the extinction coefficients, we used the OPTOOL package [50], which provides direct absorption and scattering coefficients ( and in units of cm2/g) for different types of materials and grain sizes. Previous studies based on reflectance spectra between 1998 and 2013 proposed an icy ring system around Chariklo and found that it may contain 20% water ice and 80% other materials, such as silicates, organics and small quantities of carbon [19], which motivated our choice of materials to investigate: water ice, carbon and silicates. OPTOOL offers both amorphous and crystalline options for most materials, and we found no significant differences between these options for our measurements. Therefore, we used the amorphous option in our analysis. We tested different kinds of silicates, such as pyroxene with various magnesium contents, but found minimal impact on the solutions. Therefore, we used a pyroxene model with 100% magnesium content, referred to as silicates hereafter.
In general, although is a more observationally constrained value, compositional information is carried by the opacity estimates. Model calculations (see Equation 9) can provide only optical depth values for direct comparison. However, in , this information is lost when it is combined with the radial width of the measurements. To compare the observed occultation opacities with our models, we consider the spectral energy distribution of the occulted star and each filter’s transmission curves . For a specific filter , the ratio of occulted and non-occulted photon counts is given by:
@@
First, we used single-grain models with sizes between 0.2 and 10 m. Sizes larger than 10 m do not show significant differences in extinction properties at different wavelengths in the visible to near-infrared range. We note that for planetary rings, we expect the grain size range to show a power-law distribution with the exponent around three [51]. However, with the accuracy of the occultation measurements available, our primary goal was to identify the most dominant grain sizes. In each case, we determined the best-fit models using least squares minimisation. The values were calculated as:
@@
where = with being the normal measured opacity, and is its uncertainty in the different filters. After identifying the best-fit grain sizes, we mixed the single-grain solutions using two approaches: (1) mixing the same material with two-grain sizes close to each other and (2) mixing two materials with their preferred grain sizes. The cases with the lowest values are considered as our compositional solutions for the rings.
Even though there have been a large number of occultation measurements since 2013, only a few sites could capture the very narrow C2R ring. With their remarkable sensitivity and sampling abilities, the JWST measurements provided a unique opportunity to enrich the existing observations of the secondary ring C2R. Extended Data Tables shows all the available measurements from 2013 on the C2R ring, where the JWST ones are the most well-sampled of all those measurements. Some of the earlier C2R measurements contain only one or two data points, resulting in an unresolved detection of the ring. Therefore, it is essential to have a good understanding of the , and radial width values from previous observations and their associated uncertainties in the probability space.
The most important of all is the VLT measurement in 2017 using the band, which samples the same, albeit smaller, piece of the JWST F150W2 filter range. Panel (a) in Extended Data Figures presents a comparative analysis of Chariklo’s ring system as observed with the VLT in 2017 (at 2.1 m, very close to the JWST F150W2 band) and with JWST in 2022. The radial brightness profiles reveal consistent locations and morphologies of the C1R and C2R ring components across both epochs and instruments, with only minor differences attributable to changes in viewing geometry and filter transmission curves. The JWST observations in the F150W2 and F322W2 filters provide higher signal-to-noise ratios, allowing finer sampling of the ring structure and sharper detection of the main features compared to ground-based data. Notably, however, while a stronger detection of C2R was expected with JWST, the opposite trend is observed: the C2R signal appears significantly weaker, especially in the F322W2 filter, where the z-score of the 1 and 2 contacts reaches only ~0.31 and ~0.23, respectively. Even in F150W2, where the JWST detection is somewhat clearer (~1.70 and ~1.73), it remains below the VLT levels (~4.42 and ~3.36). Nevertheless, the previously published opacity values for this observation appear asymmetrical at first glance. The 1 contact has a low opacity and well-constrained error margin (), which aligns well with the JWST measurements. However, the second contact appears completely different, with considerable uncertainty () -see Extended Data Tables. Figure Extended Data Figures Panel (b) shows the actual distributions of the possible apparent opacity values of this VLT occultation. Below the red lines, the 1- and 3- regions can be found. This is a typical result of fitting the occultation light curves using a Monte Carlo approach with different possible models and evaluating them. For the 1 contact, we can confirm that the value is fairly constrained. Still, for the second one, even though there is (numerically) a best-fitted value (), there is little to no difference between the values within a wide range (which resulted in error bars). This clearly demonstrates that, due to the low resolution of the VLT measurement, we cannot determine a single value for , but rather a range that, as with the 1 contact, overlaps with the JWST F150W2 region.
Because the previously published individual opacity and width values may not accurately represent the most likely values within their given error bars, we used a probabilistic approach to examine the typical radial width values of the C2R ring. To identify the most representative radial widths () derived from the occultation light curve fits, we performed a kernel density estimation (KDE) using all valid C2R measurements. The resulting probability density function, shown in Figure Extended Data Figures, exhibits two distinct peaks that characterise the most probable width regimes ().
These KDE-derived peaks guided the recalculation of the normal optical depth , using the relation:
where is the equivalent width obtained from the light curve fitting.
When the measured of an observation overlapped with a KDE peak, that value was directly used in Equation 12. In contrast, if the measured did not intersect any peak, a constrained search within its asymmetric uncertainty bounds was performed to select the value closest to the nearest KDE maximum, while ensuring it remained within the original asymmetric limits. From the set of candidate values yielding within the observational error limits, the extremal value (either minimum or maximum) was selected based on the direction of the nearest peak. This procedure was applied, for instance, to observations such as OPD (both contacts), Springbok (second contact), and Gifberg (second contact), which do not directly overlap with either KDE peak. This approach avoids arbitrary resampling and preserves both physical consistency and fidelity to observational uncertainties.
By applying this method across all measurements, we derived new, more probable normal optical depth values, , constrained strictly within the original asymmetric uncertainties. The resulting distribution, shown in Figure 3, illustrates the original uncertainty ranges along with the resampled values (Extended Data Tables). The KDE-based adjustment significantly reduces the spread of the data points, highlighting consistent trends that were previously obscured. For instance, the two VLT observations, which originally showed substantial differences in (see Extended Data Tables), now exhibit closer agreement once KDE-informed widths are applied. However, the second, much larger VLT measurement also covers a wider potential opacity range with an overlap with the first, well-constrained value, which also agrees well with the JWST F150W2 observation. Our analyses indicate a reasonable likelihood that the two 2017 VLT measurements exhibit similar opacities to the 2022 JWST measurements, suggesting a silicate-like composition consisting of small grains. Considering also the Danish observations, the preferred silicate solutions for the C2R ring are between 0.2 – 0.5 m. The two Danish and the two JWST observations are crucial because, for each instrument, the observations in two filters were parallel in time and sampled the same part of the ring. Therefore, they are not sensitive to possible inhomogeneities in the ring.
The bottom Panel of Figure 3 presents the comparison between the resampled (and original) observations and the models. The normal opacities transformed into relative intensities as , showing the fractional dimming of the occulted star, which now can be directly compared to the models. For the non-detection in the F322W2 filter, we indicated a violet range set to a relative intensity of 1.0, where the width of the region represents the uncertainty of the non-detection – larger signals should have been detected. We created pyroxene models with a magnesium content of 100% using 0.2, 0.3, 0.5, 1.0 and 3.0 m grain sizes, marked as red lines in the bottom panel of Figure 3. These silicate models can capture the general trend of the C2R opacities, which suggests that opacities in the visible wavelength are more prominent than the near-infrared ones. This decreasing signal strength aligns well with the fact that the smaller the grain size of the silicate models, the greater their observable opacity in the visible range. Therefore, we suggest a compatible dusty ring model for C2R dominated by silicate grains in the range of m. We also note that the optical properties of the pyroxene models are also suitable for explaining an asymmetric detection in the two JWST filters, as based on these models, we expect a higher signal in the F150W2 compared to F322W2. Since the JWST measurements in these filters were taken at the same time, sampling the same ring segment, such a silicate-dominated dusty ring may also contribute to an unexpected difference between the two filters. All these models show that in the F322W2 band, no significant signal can be detected, if grain sizes are kept below 3 m, as represented by the red curves confined to the violet-coloured region.
The sky blue lines correspond to the water models with grain sizes of 0.5, 1.0 and 4.0 m, which display a distinctive signature feature at around 3 m. This particular feature is revisited in Figure 4, where we mainly focus on the JWST filter ranges. The middle panel of Figure 4 clearly shows that the smaller the grain size, the larger the feature itself. For grains larger than 2-5 in size, the curves flatten, distributing nearly equal contribution among all wavelengths. Fortunately, the wavelength range of this water signature is entirely covered by the F322W2 passband, offering a natural selection between the suggested models in this scenario. Taking the transmission profiles of each filter into account (top Panel in Figure 4), water ice grains of m would make a notable contribution in the F322W2 filter, relative to the F150W2 one. Pure water ice rings, similar to those around Saturn [30], cannot explain the measurements for either sub-micron or larger particles. Although our numerically best fit model for all the data is clearly the pure pyroxene one (see Extended Data Figures), the finite resolution of the measurements leaves room for other, similarly suitable mixture models. Therefore, the presence of a certain amount of water ice remains compatible with the measurements. However, in this purely wavelength-dependent scenario, the exact amount of water ice in C2R heavily depends on the chosen best-fit silicate model and the assumed water grain sizes. The bottom panel of Figure 4 exhibits silicate models with grain sizes of m, combined with different amounts of m, (solid) or m (dotted) ice grains. Besides the exact differences, what is the same in all models is that water content above 10-30% would lead to a decrease in the observed visual-infrared opacity discrepancy and a significant increase in the contribution of the F322W2 filter, which would then have been observed along with the marginal detection in F150W2. These estimates of water content are also very similar to those previously suggested for the C1R ring [19].
@@
To investigate the wavelength-dependent opacity of C1R, we performed a similar, simple radiative transfer calculation assuming a homogeneous layer for each occultation event, composed of a mixture of two materials with different compositions and grain sizes. The pre-JWST data, including the 2017 VLT/Ks measurement at , can be well fitted with a mixture of silicates (olivine or pyroxene) with grain sizes and water ice with grain sizes between 1–10 (blue curve in Figure 5). This fit implies nearly constant opacity across the near-infrared, with increased opacity at shorter visible wavelengths. Comparable fits can also be achieved using mixtures of silicate grains with two distinct size populations, similar to one above. Thus, these data do not strongly constrain the material composition of C1R. However, grains larger than would produce nearly wavelength-independent opacity over the full range considered, regardless of composition, suggesting that the grains responsible for most of the extinction in C1R are likely smaller than this threshold. When the JWST data are included, however, no combination of two materials and grain sizes provides a satisfactory fit across all datasets—particularly the JWST F150W2, F322W2, and VLT/Ks bands. The best-fit models in this case (red curve in Figure 5) require predominantly silicates with 1–10 grain sizes, but the overall fit quality remains poor. This strongly suggests that the observed opacity variations cannot be explained solely by wavelength dependence. For comparison, spectroscopy of the unresolved Chariklo system indicates a composition of 20% water ice mixed with darker materials, probably silicates [19].

### v2

@@ 1.1 Predicting and observing stellar occultations from space
In June 2020, the first-ever predicted stellar occultation observation using a space telescope was obtained with the Characterising ExOPlanets Satellite (CHEOPS) [Morgado2022]. This observation demonstrated the feasibility of predicting and observing stellar occultations by distant solar system minor bodies using space telescopes: these facilities are unaffected by weather and can extend the range of observations beyond Earth’s surface. The use of space telescopes also facilitates the detection of material surrounding small bodies, enabling photometric accuracy comparable to that achieved by ground-based telescopes with much larger apertures. We also have access to stellar occultations of fainter stars than we do from the ground, thereby multiplying the number of possible events.
@@
JWST is orbiting close to the second Sun-Earth Lagrange point (L2) [Gardner2023, McElwain2023]. L2 is an unstable equilibrium point, meaning constant manoeuvres must be performed to keep the telescope in orbit around it [Menzel2023]: these are usually performed 2–3 times a month. Consequently, the predicted orbital path of JWST needs to be frequently updated (J. Giorgini, personal communication, June 22, 2022). For all the above reasons, and given the complex motion of the space telescope around L2, predicting stellar occultations by distant minor bodies visible from JWST is a very challenging task. Extended Data Figures shows the difference between two consecutive releases of the JWST ephemeris, obtained on September 27, 2022. In just under six months from the initial epoch, the difference between the two releases reached approximately 10,000 km, exceeding the Earth’s radius. In the absence of formal positional uncertainty, the plot illustrates the trustworthiness of a stellar occultation prediction using JWST far from the initial epoch.
@@
To predict stellar occultations that JWST could potentially observe, we used the Stellar Occultation Reduction and Analysis (SORA) software package [GomesJunior2022] combined with the Gaia DR3 catalogue [Gaia2016, Gaia2023] and the NIMA ephemeris [Desmars2015] for the solar system bodies. The strategy we developed to run the predictions is as follows. We selected the 40 largest TNOs/Centaurs (36 TNOs and 4 Centaurs), then ran predictions for these 40 targets every week for Gaia DR3 stars up to a G-magnitude of 20. This re-running of the predictions was necessary because of the station-keeping manoeuvres applied to JWST, which changed its position (ephemeris). Each week, we reviewed the updated predictions to identify favourable events, refined them with new ground-based astrometric data, and reran the predictions for those objects. Finally, when we were confident in the prediction, we activated the target of opportunity (ToO) to observe the occultation from JWST. Since this was a non-disruptive ToO, it had to be activated at least 14 days before the event.
@@
Using the described strategy, we identified the occultation by the ringed Centaur Chariklo on 18 October 2022 (UT), with a low relative velocity (2.5 km/s) and a star bright enough to be observed by JWST (K = 15.7 mag) as the most promising event. We first identified this occultation on 19 August 2022 (blue line in Extended Data Figures). Then we reran the prediction each week to account for changes in JWST’s orbit caused by station-keeping manoeuvres. Our most recent prediction was made on the day of the occultation, just a few hours before the event (green line in Extended Data Figures). In just about two months, the path of the star behind Chariklo, as seen by JWST, moved 110 km, enough to miss the occultation by the main body. Furthermore, the instant of occultation also shifted by 1.5 minutes.
@@ 1.2 Observations
We used the James Webb Space Telescope (JWST) to observe a stellar occultation by Chariklo’s rings, marking the first time such an event was both predicted and successfully observed by JWST. It also represents the first stellar occultation by a small Solar System body ever recorded from beyond Earth’s orbit, with the only comparable case being a previous Quaoar occultation observed by the Earth-orbiting CHEOPS satellite. This observation was both scientifically valuable and technically demanding. It was conducted under JWST’s Solar System Guaranteed Time Observations (GTO) program #1271, ToO TNOs: Unveiling the Kuiper belt by stellar occultations (PI: Pablo Santos-Sanz) [SantosSanz2017].
@@ 1.3 Photometry: occultation light curves
Fluxes from Chariklo and the target star were also measured separately in both filters, using images acquired before the occultation when the two sources were spatially resolved (Extended Data Figures, left Panel). These measurements were used to calibrate the relative contributions of Chariklo and the star to the combined flux, a crucial step in deriving the ring parameters discussed in Section 1.4. The fractional flux contribution of Chariklo (relative to the combined Chariklo + star flux) was 0.2780 for F150W2 and 0.2662 for F322W2.
@@ 1.4 Analysis of the rings
In the obtained light curves, the main, densest, and most opaque ring of Chariklo is clearly visible. This ring is called C1R, following the original notation [Braga-Ribas2014]. The C1R occultation light curve displays prominent diffraction spike features immediately before and after the ingress and egress, respectively. These spikes provide strong evidence that the ring has sharp edges and is tightly confined, similar to the rings of Uranus [Elliot1977, Elliot1987]. However, the dynamical mechanism responsible for this confinement is still unknown and is currently under discussion [Sicardy2019, Sicardy2020a, GiuliattiWinter2023, Salo2024].
The other ring (referred to as C2R) is surprisingly not clearly detected in the F150W2 or F322W2 light curves, as shown in Extended Data Figures. Since a ring partially occults the starlight, we fit an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction, finite bandwidth, exposure time, and the Gaia DR3 stellar diameter. We use the procedures within the SORA package [GomesJunior2022]. From this analysis, we obtained the time when the star disappeared and reappeared behind Chariklo’s rings ( and ) and the observed apparent ring opacity (), which measures the fractional drop of stellar flux.
Knowing the reconstructed JWST position at the time of the occultation, we were able to propagate the instants of disappearance and reappearance of the star into the tangent plane and determine two-dimensional coordinates (, ) of the rings’ positions relative to Chariklo’s expected centre. This projection was executed using the SORA package, and its result is shown in Figure 1. Assuming Chariklo’s parameters (body and rings) as published by [Morgado2021], we fitted the centre position of Chariklo as viewed by JWST at the mid-point of the occultation (18 October 2022 at 13:37:42.18 UT), which resulted in
That meant an appulse with the closest approach of 132.2 km, in contrast to Chariklo’s volume-equivalent radius of 124.8 km [Morgado2021].
Initially, the parameters of both rings (C1R and C2R) were obtained in the sky plane simultaneously, just before and after the closest approach. We used the formalism presented by [Berard2017] to project these parameters to the ring plane so we could obtain the relevant parameters, such as the ring radial width (), the normal opacity (), and the normal optical depth ().
@@
alternatively, if we assume a poly layer ring, is defined as [Elliot1984]
where is the ring opening angle, and (resp. ) corresponds to a pole-on (resp. edge-on) viewing. On the other hand, considering that the Airy scale is larger than the width of the ring as seen in the sky plane [Boissel2014], a factor is needed to compute the fraction of lost light; in that case
@@
The integrals of and over the radial width () of the ring define the equivalent width () and the equivalent depth () of the profile. These values ( and ) are proportional to the amount of material present in the profiles for each case [Berard2017]. The ring parameters are obtained in Table 1. We highlight that the detection of C2R is 1 significant in the F150W2 data and cannot be detected in the F322W2 data.
@@ 1.5 Sampling a ring with azimuthally variable structure
In the case of a ring with an azimuthally variable structure, we may observe different normal opacities () or equivalent with () values at the 1 and 2 contacts during an occultation event, as well as variations between multiple occultations. If the ring opacity is assumed to be wavelength-independent, the observed decrease/increase in opacity and/or in during the 2022 JWST events—relative to previous occultations—could be explained by the possibility that both ingress and egress sampled low/high-density segments of the ring. To estimate the likelihood of such an occurrence, we performed a simple Monte Carlo simulation. We modelled an azimuthally variable ring, with the material distribution given by or , where is the azimuthal angle and is a small integer. Random chords drawn through this ring yield a distribution of ‘observed’ or values.
In the real occultation measurements of the C2R ring, we obtain in seven of the occultations (see Extended Data Tables), with significantly lower values observed only during the JWST event. For simplicity, we assume that the minimum of along the ring is zero, although a constant positive offset may exist. Therefore, the probability of randomly sampling a pair of such low values in just one of eight events, as calculated from the simulation, represents an upper limit, with a probability of obtained in our simulation. Similarly, we obtain a probability of p 0.004, assuming that we sample the ring near the density maxima in the JWST measurements and below these values in the pre-JWST cases (Extended Data Tables). While azimuthal fragmentation is plausible—especially given the existence of fragmented rings in the trans-Neptunian region, e.g. in the case of Quaoar [Morgado2022, Proudfoot2025]—these very low probabilities suggest that the observed JWST opacity drop (C2R) or increase (C1R) is unlikely to be solely due to this effect. The results are independent of the parameter for 3. Here, we assumed that the azimuthal structures of the two rings are independent.
@@ 1.6 Spreading timescale of Chariklo’s C2R Ring
We compute the characteristic spreading timescale of Chariklo’s second ring (C2R) using the standard expression for a narrow, self-gravitating ring in resonance [DePater2018]:
@@
We adopted ellipsoidal semi-axes from Morgado et al. (2021)[Morgado2021]: km, km and km and a mean bulk density of .
@@
To account for the uncertainty in the radial width, we uniformly sampled over the interval defined by the minimum and maximum C2R widths measured across all occultations, [0.10, 3.72] km (see Extended Data Tables). A Monte Carlo simulation with realisations yields a median spreading timescale of s (0.34 yr), with a 1 confidence interval ranging from s (0.21 yr) to s (0.96 yr).
@@ 1.7 Radiative Transfer Modeling
To explore whether a potential wavelength-dependent effect is responsible for the observed changes in the measured opacities for C1R, we conducted a radiative transfer analysis, assuming that the ring did not change during the observation period (2013-2022) and that all observational chords sampled the same amount of material, after correcting for the opening angle change due to Chariklo’s orbital motion around the Sun. We used a slab-like, two-component model, where we mixed materials to match their common optical properties with the measured occultation normal opacities () or, equivalently, their normal optical depths (). In our model, the optical depth is obtained as the sum of the optical depths of two materials (two different grain sizes and/or compositions):
@@
where is the slab’s length (the ring’s thickness). A single-component model is obtained if is chosen. To calculate the extinction coefficients, we used the OPTOOL package [optool], which provides direct absorption and scattering coefficients ( and in units of cm2/g) for different types of materials and grain sizes. Previous studies based on reflectance spectra between 1998 and 2013 proposed an icy ring system around Chariklo and found that it may contain 20% water ice and 80% other materials, such as silicates, organics and small quantities of carbon [Duffard2014], which motivated our choice of materials to investigate: water ice, carbon and silicates. OPTOOL offers both amorphous and crystalline options for most materials, and we found no significant differences between them in our measurements. Therefore, we used the amorphous option in our analysis. We tested different types of silicates, including pyroxene with varying magnesium content, but found minimal impact on the solutions. Therefore, we used a pyroxene model with 100% magnesium content, referred to as silicates hereafter.
In general, although is more observationally constrained, compositional information is carried by the opacity estimates. Model calculations (see Equation 9) can provide only optical depth values for direct comparison. However, in , this information is lost when it is combined with the radial width of the measurements. To compare the observed occultation opacities with our models, we consider the spectral energy distribution of the occulted star and each filter’s transmission curves . For a specific filter , the ratio of occulted and non-occulted photon counts is given by:
@@
First, we used single-grain models with sizes ranging from 0.2 to 10 m. Sizes larger than 10 m do not show significant differences in extinction properties at different wavelengths in the visible to near-infrared range. We note that for planetary rings, we expect the grain size range to show a power-law distribution with the exponent around three [Brilliantov2015]. However, given the available accuracy of the occultation measurements, our primary goal was to identify the dominant grain sizes. In each case, we determined the best-fit models using least squares minimisation. The values were calculated as:
@@
where = denotes the normal fractional transmission, with the measured normal opacity, and its associated uncertainty in the different filters. After identifying the best-fit grain sizes, we mixed the single-grain solutions using two approaches: (1) mixing the same material with two-grain sizes close to each other and (2) mixing two materials with their preferred grain sizes. The cases with the lowest values are considered as our compositional solutions for the rings.
@@
Here, we perform a thorough analysis of the compatibility of dusty ring models for C1R, assuming a homogeneous layer for each occultation event, composed of a mixture of two materials with different compositions and grain sizes. The pre-JWST data, including the 2017 VLT/Ks measurement at , can be well fitted with a mixture of silicates (olivine or pyroxene) with grain sizes and water ice with grain sizes between 1–10 (blue curve in Figure 3). This fit implies nearly constant opacity across the near-infrared, with increased opacity at shorter visible wavelengths. Comparable fits can also be achieved using mixtures of silicate grains with two distinct size populations, similar to one above. Thus, these data do not strongly constrain the material composition of C1R. However, grains larger than would produce nearly wavelength-independent opacity over the full range considered, regardless of composition, suggesting that the grains responsible for most of the extinction in C1R are likely smaller than this threshold. When JWST data are included, however, no combination of two materials and grain sizes provides a satisfactory fit across all datasets—particularly the JWST F150W2, F322W2, and VLT/Ks bands. The best-fit models in this case (red curve in Figure 3) require predominantly silicates with 1–10 grain sizes, but the overall fit quality remains poor. This strongly suggests that the observed opacity variations cannot be explained solely by wavelength dependence. For comparison, spectroscopy of the unresolved Chariklo system indicates a composition of 20% water ice mixed with darker materials, probably silicates [Duffard2014].